# Pólya’s conjecture up to $\epsilon$-loss and quantitative estimates for the remainder of Weyl’s law 00footnotetext: 2020 Mathematics Subject Classification. Primary 35P15; Secondary 35P20, 42B37. Key words and phrases: Dirichlet eigenvalue, Weyl law, Pólya conjecture

### v2

@@ 1.4 Our method and plan of the paper
The main strategy to prove Theorem 1.5, Theorem 1.3 and Theorem 1.9 is as following.
We first use the explicit formula of Riesz means of Dirichlet eigenvalues in ${\mathbb{R}}$ (cf. Lemma 2.1), following the inspiring observation by Laptev [26], to give a refined estimates on product domains $\Omega_{1}\times\Omega_{2}$ (see Theorem 2.3 below). This allows us to conclude via monotonicity of Dirichlet eigenvalues (cf. [18]) that for any open domains $\Omega$ contained in $\Omega_{1}\times\Omega_{2}$, the first $k_{0}$ Dirichlet eigenvalues satisfies Pólya’s conjecture (cf. Corollary 2.4).
We then establish a quantitative lower bound for the counting function (upper bound for the Dirichlet eigenvalue) on cubes (see Lemma 3.3 and Lemma 3.6 below), which together with our refinements on product domains and monotonicity of Dirichlet eigenvalues to conclude the proof Theorem 1.9. In fact, in higher dimensions, $n\geq 3$, the rectangle $\mathscr{R}=\mathscr{R}_{n-1}\times I$ can be replaced some more general product $\Omega\times I$, as soon as Pólya’s conjecture holds on $\Omega$, see Theorem 3.7 below.
Although only a part of discussions in Sections 2 and 3 will be sufficient for the proofs of Theorem 1.3 and Theorem 1.5, the extended discussions are of independent interest and they have other uses. For Theorem 1.5 we shall apply the Whitney decomposition of a Lipschitz domain $\Omega$, and using the lower bound for the counting function on cubes (see Lemma 3.3 and Lemma 3.6 below) along with an optimization to deduce a lower bound for the counting function $\mathcal{N}_{\Omega}^{D}(\lambda)$ on $\Omega$. Finally, applying our refinements on product domains (see Theorem 2.3 below) and monotonicity of Dirichlet eigenvalues, together with the lower bound of $\mathcal{N}_{\mathscr{R}\setminus\overline{\Omega}}^{D}(\lambda)$ to conclude the proof. Theorem 1.3 then follows almost immediately from Theorem 1.5.
Finally, let us point it out, our approach depends on a modified min-max argument, which is related to but different from the original ones by Courant [5] and Courant-Hilbert [6]. By our choose of minimal Admissible rectangles, we do not have to take care of the term in Courant [5] and Courant-Hilbert [6] for cubes which intersect the boundary of $\Omega$. The latter is a difficult issue for Courant [5] and Courant-Hilbert [6].
@@
We shall employ Lemma 3.3 and Lemma 3.6 from Section 3 and Theorem 2.3 from Section 2 to prove the uniform estimates on Lipschitz sets, Theorem 1.5 and its corollary, Theorem 1.3. We shall also provide a uniform estimate on general open domains, see Theorem 4.2 below.
@@
Added in remark. After the first version of this article was posted on arXiv, we are kindly informed by Frank and Larson, that a similar version to our uniform remainder estimate for Weyl’s law (Theorem 1.5) already appears in [9, Theorem 3.14] and [12, Proposition 3.6], which both show that
(1.13) $\left|\mathcal{N}^{D}_{\Omega}(\lambda)-\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\right|$ $\leq C(n)\Theta_{\Omega}\lambda^{\frac{n-1}{2}}\left(1+\log_{+}(r_{in}(\Omega)\sqrt{\lambda})\right),$
where $C(n)$ is an implicit constant depending only on the dimension $n$,
$\Theta_{\Omega}:=\sup_{\delta>0}\delta^{-1}|\left\{x\in\Omega:\,{\mathop{\mathrm{missing}}{\,dist\,}}(x,\partial\Omega)<\delta\right\}|<\infty,$
$\Theta_{\Omega}$ plays the same role of $C_{Lip}(\Omega)|\partial\Omega|$ in our estimate. In particular, it appears from the proof of [9, Theorem 3.14] that the above constant $C(n)$ can also be calculated. The proof of [12, Proposition 3.6] uses Tauberian theory, while the proof of [9, Theorem 3.14] uses Courant’s method of Dirichlet-Neumann bracketing and some refined estimates for cubes intersecting the boundary.
Our proof for estimates of the reminder of Weyl’s law (Theorem 1.5) uses the minimal Admissible rectangle (see Theorem 2.3 below for refined eigenvalue estimates on product domains), a subtracting argument that allows us to avoid the cubes intersecting the boundary, and then lower bound for the counting function of Dirichlet eigenvalues. Moreover, our method has an advantage that gives many “irregular shaped domains” that verify the Pólya conjecture (see Theorem 1.9, Remark 1.11 and Theorem 3.7 below).
2 Refined eigenvalue estimates on product domains
In this section, we prove refined eigenvalue estimates on product domains, Theorem 2.3 and its corollary, Corollary 2.4. Recall that, if an open domain $\Omega$ is in the real line, then its eigenvalues have explicit lower bounds, which satisfy for $k\geq 1$
@@ 2 Refined eigenvalue estimates on product domains
(2.1) $\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})$ $\leq\frac{2|\Omega|}{3\pi}\lambda^{3/2}\min\left\{\left(1-\frac{3\pi}{16|\Omega|\sqrt{\lambda}}\right)_{+},\,\left(1-\left(\frac{\lfloor\frac{|\Omega|\sqrt{\lambda}}{\pi}\rfloor}{\frac{|\Omega|\sqrt{\lambda}}{\pi}}\right)^{2}\frac{3\pi}{4|\Omega|\sqrt{\lambda}}\right)\right\},$
@@
(2.2) $\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})^{1/2}$ $\leq\frac{|\Omega|}{4}\lambda-\frac{\sqrt{\lambda}}{2}+\frac{\sqrt{6}\pi}{9|\Omega|}\sqrt{\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}\leq\frac{|\Omega|}{4}\lambda\left(1-\frac{2}{3}\frac{1}{|\Omega|\sqrt{\lambda}}\right).$
@@
(2.3) $\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})^{1/2}$ $\geq\frac{|\Omega|}{4}\lambda\left(1-\frac{2\pi}{|\Omega|\sqrt{\lambda}}+\frac{\pi^{2}}{|\Omega|^{2}\lambda}\right).$
@@
Let us consider the ball $B(0,\frac{|\Omega|}{\pi}\sqrt{\lambda})$ in the first quadrant in ${\mathbb{R}}^{2}$. For each $1\leq k\leq k_{\lambda}=\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor$, $(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}$ equals the area of the rectangle $[k-1,k]\times[0,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}]$, which is contained in $B(0,\frac{|\Omega|}{\pi}\sqrt{\lambda})$ in the first quadrant in ${\mathbb{R}}^{2}$. Moreover, for each $k$, there is at least a triangle with vertices at $(k-1,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-(k-1)^{2})^{1/2})$, $(k-1,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2})$ and $(k,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2})$ contained in the ball. Moreover, if $k_{\lambda}\neq\frac{|\Omega|}{\pi}\sqrt{\lambda}$, there is another triangle insider the 1/4 ball with vertices at $(k_{\lambda},(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k_{\lambda}^{2})^{1/2})$, $(k_{\lambda},(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k_{\lambda}^{2})^{1/2})$ and $(\frac{|\Omega|}{\pi}\sqrt{\lambda},0)$. See Figure $2$Picture drawn by Xinyi Jiang..
@@
Note that the union of the rectangles $[k-1,k]\times[0,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}]$, $1\leq k\leq k_{\lambda}$, contains the ball $B(0,\sqrt{\frac{|\Omega|^{2}}{\pi^{2}}\lambda-2\sqrt{\frac{|\Omega|^{2}}{\pi^{2}}\lambda}+1})$ intersecting the first quadrant, since
@@
(2.4) $C_{3}(p)=\begin{cases}\frac{1}{8},&p=1,\\ \frac{1}{9},&1<p\leq 2,\\ \frac{1}{8}\left(\frac{247}{256}\right)^{p-1},&p>2.\end{cases}$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
(2.7) $C_{1}(n_{1})=\begin{cases}\frac{2}{3}&n_{1}=1,\\ \frac{3\pi}{16}&n_{1}=2,\\ \frac{\pi\mathscr{B}(\frac{n_{1}}{2}-1,2)}{6\mathscr{B}(\frac{n_{1}}{2}-1,5/2)}&n_{1}=3,4,\\ \frac{3\pi\mathscr{B}(\frac{n_{1}}{2}-1,2)}{16\mathscr{B}(\frac{n_{1}}{2}-1,5/2)}\left(\frac{247}{256}\right)^{\frac{n_{1}-2}{2}}&n_{1}\geq 5.\end{cases}$
@@
(2.8) $\lambda_{1}(\Omega)=\lambda_{1}(\Omega_{1})+\lambda_{1}(\Omega_{2})\geq\frac{(2\pi)^{2}}{|\Omega_{1}|^{2/n}\omega(n_{1})^{2/n_{1}}}+\frac{\pi^{2}}{|\Omega_{2}|^{2}},$
@@
(2.9) $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}.$
@@
Let $0\leq k\leq k_{0}$. If $\lambda_{k}(\Omega)<\lambda_{0}$, then (2) gives that for small enough $\epsilon>0$ such that $\lambda_{k}(\Omega)+\epsilon<\lambda_{0}$

### v3

@@ 1.4 Our method and plan of the paper
The main strategy of the paper is as following.
We first use the explicit formula of Riesz means of Dirichlet eigenvalues in $\mathbb{R}$} (cf. Lemma 2.1), following the inspiring observation by Laptev [26], to give a refined estimates on product domains $\Omega_{1}\times\Omega_{2}$ (see Theorem 2.3 below). Using these refined estimates on product domains together with Pólya’s method [35], we give the refined estimates on strip-tiling domains.
We then establish a quantitative lower bound for the counting function (upper bound for the Dirichlet eigenvalue) on cubes (see Lemma 3.3 and Lemma 3.6 below), which together with our refinements on strip-tiling domains and monotonicity of Dirichlet eigenvalues to conclude the proof Theorem 1.15. Moreover, in higher dimensions, $n\geq 3$, certain product domains can replace the role of strip-tiling domains, see Theorem 3.7 below.
For Theorem 1.5 we shall apply the Whitney decomposition of a Lipschitz domain $\Omega$, and using the lower bound for the counting function on cubes (see Lemma 3.3 and Lemma 3.6 below) along with an optimization to deduce a lower bound for the counting function $\mathcal{N}_{\Omega}^{D}(\lambda)$ on $\Omega$. Finally, applying our refinements on product domains (see Theorem 2.3 below) and monotonicity of Dirichlet eigenvalues, together with the lower bound of $\mathcal{N}_{\mathscr{R}\setminus\overline{\Omega}}^{D}(\lambda)$ to conclude the proof. Theorem 1.3 then follows almost immediately from Theorem 1.5.
Finally, let us point it out, our approach depends on a modified min-max argument, which is related to but different from the original ones by Courant [5] and Courant-Hilbert [6]. By our choose of minimal Admissible rectangles, we do not have to use Neumann eigenvalues which is crucial in Courant’s Dirichlet-Neumann bracketing method (see [5, 6, 9, 32]), and gives better constant estimates.
@@
We shall employ Lemma 3.3 and Lemma 3.6 from Section 3 and Theorem 2.3 from Section 2 to prove the uniform estimates on Lipschitz sets, Theorem 1.5 and its corollary, Theorem 1.3.
@@
2 Refined eigenvalue estimates
In this section, we prove refined eigenvalue estimates on product domains and strip-tiling domains.
2.1 Product domains
We first prove Theorem 2.3 and its corollary, Corollary 2.4. Recall that, if an open domain $\Omega$ is in the real line, then its eigenvalues have explicit lower bounds, which satisfy for $k\geq 1$
@@ 2.1 Product domains
$\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})$ $\leq\frac{2|\Omega|}{3\pi}\lambda^{3/2}\min\left\{\left(1-\frac{3\pi}{16|\Omega|\sqrt{\lambda}}\right)_{+},\,\left(1-\left(\frac{\lfloor\frac{|\Omega|\sqrt{\lambda}}{\pi}\rfloor}{\frac{|\Omega|\sqrt{\lambda}}{\pi}}\right)^{2}\frac{3\pi}{4|\Omega|\sqrt{\lambda}}\right)\right\},$ (2.1)
@@
$\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})^{1/2}$ $\leq\frac{|\Omega|}{4}\lambda-\frac{\sqrt{\lambda}}{2}+\frac{\sqrt{6}\pi}{9|\Omega|}\sqrt{\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}\leq\frac{|\Omega|}{4}\lambda\left(1-\frac{2}{3}\frac{1}{|\Omega|\sqrt{\lambda}}\right).$ (2.2)
@@
$\sum_{k:\,\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k})^{1/2}$ $\geq\frac{|\Omega|}{4}\lambda\left(1-\frac{2\pi}{|\Omega|\sqrt{\lambda}}+\frac{\pi^{2}}{|\Omega|^{2}\lambda}\right).$ (2.3)
@@
Let us consider the ball $B(0,\frac{|\Omega|}{\pi}\sqrt{\lambda})$ in the first quadrant in ${\mathbb{R}}^{2}$. For each $1\leq k\leq k_{\lambda}=\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor$, $(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}$ equals the area of the rectangle $k-1,k]\times[0,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}]$, which is contained in $B(0,\frac{|\Omega|}{\pi}\sqrt{\lambda})$ in the first quadrant in ${\mathbb{R}}^{2}$. Moreover, for each $k$, there is at least a triangle with vertices at $(k-1,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-(k-1)^{2})^{1/2})$, $(k-1,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2})$ and $(k,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2})$ contained in the ball. Moreover, if $k_{\lambda}\neq\frac{|\Omega|}{\pi}\sqrt{\lambda}$, there is another triangle insider the 1/4 ball with vertices at $(k_{\lambda},(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k_{\lambda}^{2})^{1/2})$, $(k_{\lambda},(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k_{\lambda}^{2})^{1/2})$ and $(\frac{|\Omega|}{\pi}\sqrt{\lambda},0)$. See Figure 3222Picture drawn by Xinyi Jiang..
@@
Note that the union of the rectangles $k-1,k]\times[0,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}]$, $1\leq k\leq k_{\lambda}$, contains the ball $B(0,\sqrt{\frac{|\Omega|^{2}}{\pi^{2}}\lambda-2\sqrt{\frac{|\Omega|^{2}}{\pi^{2}}\lambda}+1})$ intersecting the first quadrant, since
@@
$C_{3}(p)=\begin{cases}\frac{1}{8},&p=1,\\ \frac{1}{9},&1<p\leq 2,\\ \frac{1}{8}\left(\frac{247}{256}\right)^{p-1},&p>2.\end{cases}$ (2.4)
@@
$\sum_{k:\,\lambda_{k}<\lambda}(\lambda-\lambda_{k})^{p}$ $\leq\mathscr{B}(p-1,2)^{-1}\sum_{k}\int_{0}^{\infty}v^{p-2}(\lambda-\lambda_{k}-v)_{+}\,dv$
$\leq\frac{2|\Omega|}{3\pi\mathscr{B}(p-1,2)}\int_{0}^{\infty}v^{p-2}(\lambda-v)_{+}^{3/2}\left(1-\frac{3\pi}{16|\Omega|\sqrt{(\lambda-v)_{+}}}\right)_{+}\,dv$
$\leq\frac{2|\Omega|}{3\pi\mathscr{B}(p-1,2)}\lambda^{p+\frac{1}{2}}\mathscr{B}(p-1,5/2)-\frac{1}{8\mathscr{B}(p-1,2)}\int_{0}^{\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}}}v^{p-2}(\lambda-v)_{+}\,dv.$ (2.5)
@@
$\frac{1}{8\mathscr{B}(p-1,2)}\int_{0}^{\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}}}v^{p-2}(\lambda-v)_{+}\,dv\geq\frac{\lambda^{p}}{8\mathscr{B}(p-1,2)}\frac{249}{256}\mathscr{B}(p-1,2)\geq\frac{\lambda^{p}}{9}.$ (2.6)
@@
$\int_{\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}}}^{\lambda}v^{p-2}(\lambda-v)_{+}\,dv$ $=\frac{\lambda^{p}}{p-1}-\frac{\lambda^{p}}{p}-\left(\frac{\lambda(\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}})^{p-1}}{p-1}-\frac{(\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}})^{p}}{p}\right)$
@@
$\frac{1}{8\mathscr{B}(p-1,2)}\int_{0}^{\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}}}v^{p-2}(\lambda-v)_{+}\,dv$
$=\frac{1}{8\mathscr{B}(p-1,2)}\left(\int_{0}^{\lambda}v^{p-2}(\lambda-v)_{+}\,dv-\int_{\lambda-\frac{9\pi^{2}}{256|\Omega|^{2}}}^{\lambda}v^{p-2}(\lambda-v)_{+}\,dv\right)$
@@
$C_{1}(n_{1})=\begin{cases}\frac{2}{3}&n_{1}=1,\\ \frac{3\pi}{16}&n_{1}=2,\\ \frac{\pi\mathscr{B}(\frac{n_{1}}{2}-1,2)}{6\mathscr{B}(\frac{n_{1}}{2}-1,5/2)}&n_{1}=3,4,\\ \frac{3\pi\mathscr{B}(\frac{n_{1}}{2}-1,2)}{16\mathscr{B}(\frac{n_{1}}{2}-1,5/2)}\left(\frac{247}{256}\right)^{\frac{n_{1}-2}{2}}&n_{1}\geq 5.\end{cases}$ (2.7)
@@
$\lambda_{1}(\Omega)=\lambda_{1}(\Omega_{1})+\lambda_{1}(\Omega_{2})\geq\frac{(2\pi)^{2}}{|\Omega_{1}|^{2/n}\omega(n_{1})^{2/n_{1}}}+\frac{\pi^{2}}{|\Omega_{2}|^{2}},$ (2.8)
@@
$\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}.$ (2.9)
@@
Let $0\leq k\leq k_{0}$. If $\lambda_{k}(\Omega)<\lambda_{0}$, then (2.1) gives that for small enough $\epsilon>0$ such that $\lambda_{k}(\Omega)+\epsilon<\lambda_{0}$
@@
2.2 Strip-tiling domains
We next use Theorem 2.3 together with Pólya’s original method to prove Theorem 1.10.
Proof of Theorem 1.10.
Given a strip-tiling domain $\Omega$, let us consider $\ell^{n-1}$ isometries of $\Omega$, $\{\Omega_{j}\}_{1\leq j\leq\ell^{n-1}}$ that are placed as the tiling procedure such that in the first n1n-1 directions, each line only intersects at most $\ell$ isometries of $\Omega$.
Consider the minimal Admissible rectangle $\mathscr{R}=\mathscr{R}_{n-1}\times(0,w(\Omega))$ that contains these $\ell^{n-1}$ isometries of $\Omega$.
Then we have
(i) $|\mathscr{R}_{n-1}|\geq\ell^{n-1}|S_{n}(\Omega)|$,
(ii) $|\mathscr{R}\setminus(\cup_{j=1}^{\ell^{n-1}}\overline{\Omega_{j}})|\leq C(\Omega)\ell^{n-2}$. This is true since $\Omega$ tiles the strip ${\mathbb{R}}^{n-1}\times[0,w(\Omega)]$, and the set $\mathscr{R}\setminus(\cup_{j=1}^{\ell^{n-1}}\overline{\Omega_{j}})$ locates near the 2(n1)2(n-1) faces of $\mathscr{R}$ in the first n1n-1 directions. See the following Figure 4.
Figure 4: Examples of strip-tiling domains in two dimension
By Theorem 2.3, we see that for each $k\geq 1$,
$\ell^{n-1}N^{D}_{{\Omega}}(\lambda)$ $\leq N^{D}_{\cup_{j=1}^{\ell^{n-1}}\overline{\Omega_{j}}}(\lambda)\leq N^{D}_{\mathscr{R}}(\lambda)$
$\leq\frac{|\mathscr{R}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\left(1-\frac{C_{1}(n-1)}{|w(\Omega)|\sqrt{\lambda}}\right)_{+},$
which implies that
$N^{D}_{{\Omega}}(\lambda)$ $\leq\frac{\omega(n)}{(2\pi)^{n}}\frac{|\mathscr{R}|}{\ell^{n-1}}\lambda^{n/2}\left(1-\frac{C_{1}(n-1)}{|w(\Omega)|\sqrt{\lambda}}\right)_{+}.$
Sending $\ell$ to $\infty$, using the properties (i) and (ii) of $\mathscr{R}$, we see that for each $k\geq 1$,
$k$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-C_{1}(n-1)|S_{n}(\Omega)|\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}.$
The proof is complete. ∎
Remark 2.5.
In Theorem 2.3 and Theorem 1.10, we only keep one lower order term. One can have a better estimate especially for large eigenvalues, by keeping more lower order term. We present it below in two and three dimensions.
Let $\Omega_{1}\subset{\mathbb{R}}^{n_{1}}$ and $\Omega_{2}\subset{\mathbb{R}}$ be bounded open domains, $n_{1}=1,2$. If $n_{1}=1$, then by (2.2), we have
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\frac{|\Omega_{1}|}{\pi}(\lambda-\lambda_{j}(\Omega_{2}))^{1/2}\leq\frac{|\Omega_{1}|}{\pi}\left(\frac{|\Omega_{2}|}{4}\lambda-\frac{\sqrt{\lambda}}{2}+\frac{\sqrt{6}\pi}{9|\Omega_{2}|}\sqrt{\lfloor\frac{|\Omega_{2}|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}\right)_{+}$
$\leq\frac{|\Omega_{1}||\Omega_{2}|}{4\pi}\lambda\left(1-\frac{2}{|\Omega_{2}|\sqrt{\lambda}}+\frac{4\sqrt{6}\pi}{9|\Omega_{2}|^{2}\lambda}\sqrt{\lfloor\frac{|\Omega_{2}|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}\right)_{+}.$
If $n_{1}=2$, we assume that Pólya’s conjecture holds on $\Omega_{1}$. By (2.1) we have
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\sum_{j:\,\lambda_{j}(\Omega_{2})<\lambda}\frac{|\Omega_{1}|\omega(2)}{(2\pi)^{2}}(\lambda-\lambda_{j}(\Omega_{2}))$
$\leq\frac{|\Omega_{1}||\Omega_{2}|}{6\pi^{2}}\lambda^{3/2}\left(1-\left(\frac{\lfloor\frac{|\Omega_{2}|\sqrt{\lambda}}{\pi}\rfloor}{\frac{|\Omega_{2}|\sqrt{\lambda}}{\pi}}\right)^{2}\frac{3\pi}{4|\Omega_{2}|\sqrt{\lambda}}\right)_{+}.$
Let $\Omega$ be a strip-tiling domain that can tile ${\mathbb{R}}^{n-1}\times(0,w(\Omega))$, $n=2,3$, with $S_{n}(\Omega)$ be a face in the $n$-direction. From the proof of Theorem 1.10 we deduce similar estimate as following. For the case $n=2$, it holds that
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega|}{4\pi}\lambda\left(1-\frac{2}{|w(\Omega)|\sqrt{\lambda}}+\frac{4\sqrt{6}\pi}{9|w(\Omega)|^{2}\lambda}\sqrt{\lfloor\frac{|w(\Omega)|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}\right)_{+},$
and for $n=3$, it holds that
$\mathcal{N}^{D}_{\Omega}(\lambda)$ $\leq\frac{|\Omega|}{6\pi^{2}}\lambda^{3/2}\left(1-\left(\frac{\lfloor\frac{|w(\Omega)|\sqrt{\lambda}}{\pi}\rfloor}{\frac{|w(\Omega)|\sqrt{\lambda}}{\pi}}\right)^{2}\frac{3\pi}{4|w(\Omega)|\sqrt{\lambda}}\right)_{+}.$
Similar to Corollary 2.4 we have the following perturbation result.
Corollary 2.6.
Let $\Omega$ be a strip-tiling domain that can tile ${\mathbb{R}}^{n-1}\times(0,w(\Omega))$, $n\geq 2$. Let $S_{n}(\Omega)$ be a side face of $\Omega$ in the $n$-th direction. Let $\Omega_{0}$ be an open set contained in $\Omega$, and
$k_{0}=\left\lfloor\frac{|\Omega_{0}|\omega(n)}{(2\pi)^{n}}\left[\left(1-\frac{|\Omega_{0}|}{|\Omega|}\right)^{-1}\frac{C_{1}(n-1)}{|w(\Omega)|}\right]^{n}\right\rfloor.$
$k\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}.$